# ClipRover: Zero-shot Vision-Language Exploration and Target Discovery by Mobile Robots

### v1

@@ Abstract
Vision-language navigation (VLN) has emerged as a promising paradigm, enabling mobile robots to perform zero-shot inference and execute tasks without specific pre-programming. However, current systems often separate map exploration and path planning, with exploration relying on inefficient algorithms due to limited (partially observed) environmental information. In this paper, we present a novel navigation pipeline named “ClipRover” for simultaneous exploration and target discovery in unknown environments, leveraging the capabilities of a vision-language model named CLIP. Our approach requires only monocular vision and operates without any prior map or knowledge about the target. For comprehensive evaluations, we design the functional prototype of a UGV (unmanned ground vehicle) system named “Rover Master”, a customized platform for general-purpose VLN tasks. We integrate and deploy the ClipRover pipeline on Rover Master to evaluate its throughput, obstacle avoidance capability, and trajectory performance across various real-world scenarios. Experimental results demonstrate that ClipRover consistently outperforms traditional map traversal algorithms and achieves performance comparable to path-planning methods that depend on prior map and target knowledge. Notably, ClipRover offers real-time active navigation without requiring pre-captured candidate images or pre-built node graphs, addressing key limitations of existing VLN pipelines.
@@ 1 Introduction
Autonomous exploration and target discovery is a fundamental problem in robotics, with applications ranging from search-and-rescue to environmental monitoring and surface mapping [70, 43]. In 2D exploration, notable implementations are seen in household robots, e.g., vacuum cleaning [23, 65] and warehouse robotics [16, 34]. Traditionally, their navigation strategies have relied on various forms of Bug algorithms (Bug0, Bug1, Bug2) [45, 47], which are best suited when a robot has a priori belief about the environment, and can sense obstacles locally. The principle is to use the line-of-sight gradient towards the goal by applying simple obstacle avoidance schemes. Algorithms such as random walks or predefined search patterns are also utilized to explore an unknown workspace to develop an occupancy grid map, on which standard path-planning algorithms [69, 3] can be deployed for mobile robot navigation. Tree search algorithms like Dijkstra and A* [9] have been popular as they are simple and efficient for limited onboard resources [73]. For large search spaces, sampling-based methods such as Probabilistic Roadmaps (PRM) [33] and Rapidly-exploring Random Tree (RRT) [39] and RRT* [32] are generally adopted for source-to-target path planning in real-time.
In existing unmanned ground vehicle (UGV) systems, mapping and target discovery are typically conducted as separate processes, relying on different sensors. For instance, a 2D LiDAR is commonly used for obstacle avoidance and mapping [66, 57], while a separate camera is used for scene classification and target detection [35]. In a more complicated 3D environment such as an underwater cave exploration, the robot may need to use a combination of sensors such as sonar, LiDAR, and cameras to navigate and explore the environment by generating a semantic representation of the surroundings [1, 59]. However, an initial map traversal is not feasible in partially observable and dynamic environments, requiring frequent correction or re-initialization [24, 7].
More advanced UGV systems address this by using online path planning algorithms [53, 54] with dynamic obstacles and partially observable state space. The major challenge here is to adapt the path when new obstacles intersect the trajectory and/or dynamic changes occur in the environment [13]. To this end, visually-guided robots use the semantic understanding and spatio-temporal modeling of dynamic objects in the scene [26, 50]. For visual semantic navigation, the path planner becomes an intermediate representation of a task planner [60] for application-specific higher-level tasks such as inspection [72] search-and-rescue [71], exploration [41], etc. In recent years, more interactive task planners are designed with vision-language navigation (VLN) models [25, 17] to enable human-like understanding of the robot’s tasks from natural language and visual (semantic) cues. These are particularly useful for human-in-the-loop systems [28], e.g., robots operating in warehouses, healthcare, and social settings. VLNs enable robots to interpret the context of the tasks from language description and map the corresponding visual cues to make context-aware navigation decisions. However, integrating vision-language models (VLMs) into the early stages of map exploration – remains unexplored. We hypothesize that this integration can enable robots to leverage high-level zero-shot visual information for simultaneous exploration and target discovery without a prior map.
This paper presents ClipRover, a novel framework that utilizes the spatial context awareness capabilities of general-purpose VLMs [10] to guide robotic exploration and target discovery in unknown environments. The framework adopts a modular architecture organized into three key stages: perception, correlation, and decision. This modular design enables the decomposition of complex system-level challenges into smaller, manageable sub-problems, allowing for iterative improvements to each stage. Furthermore, the framework features a highly configurable correlation middleware, offering flexibility to adapt to various tasks and environmental conditions without any prior map or knowledge about the target.
To validate the proposed framework in real-world environments, we develop a UGV platform that meets the computational demands of VLMs while offering superior maneuverability and mechanical stability during the task. We deployed the CLIP model [58, 27] in our proposed architecture and performed extensive tests in a real-world environment. Overall, we make the following contributions to this paper:
@@
    Science: We propose a novel navigation pipeline named ClipRover for simultaneous exploration and target discovery by UGVs in unknown environments. It harnesses the power of VLMs into a modular architecture, consisting of a vision-encoder as frontend, language correlation databases as middlewares, and a decision-maker as the backend.
@@
    System: We design a novel UGV platform to support the maneuverability and computational demands of VLM-based navigation systems like ClipRover. It is designed to be powerful and configurable for general-purpose robotics research; it will be open-sourced for academic use.
@@
    Integration: We develop a fully functional system integrating the proposed navigation pipeline with the UGV platform. It is then optimized for real-time performance through comprehensive benchmarks. We present the key challenges and practicalities involved in deploying such systems in real-world application scenarios.
We conducted extensive experiments to evaluate the proposed system against both map-traversal and path-finding algorithms. A performance comparison based on efficiency and success rates is shown in Fig. 2, identifying four equipotential lines: EP0-EP3. In our evaluation, we observe that algorithms that use a similar amount of information tend to reside on the same EP. A higher equipotential line represents more information (e.g., prior map of the environment, target location, additional sensors) available to the algorithm. With zero prior information of the environment (EP0), ClipRover offers performance margins comparable and often better than Bug2 (EP2-EP3). Note that Bug2 needs prior information on the target position and a precise localization to keep track of its m-line. Overall, ClipRover delivers $\mathbf{33\%}$ more efficient source-to-target paths compared to Bug2 and achieves a $\mathbf{10\%}$ higher success than the Bug1 algorithm. Importantly, none of the evaluated planner and traversal methods have a semantic understanding of the scene, limiting their ability to adapt to dynamic environments.
With vision-language integration, SOTA systems such as Clip-Nav [14], Clip of Wheels [17], and GCN [36] - demonstrate superior navigation performance while providing semantic scene understanding. However, these systems rely on pre-captured images of candidate scenes, target locations, and pre-built knowledge graphs of the navigable space [11]. Additionally, many are tested primarily in simulation or overly simplistic environments. Other contemporary methods, such as SEEK [19], require high-sensing modalities involving multiple cameras, LiDAR, and onboard SLAM pipelines for navigation. ClipRover overcomes these limitations by enabling real-time active navigation using only a monocular camera, without relying on any prior information or pre-built dependencies.
Figure 2: Performance of ClipRover compared to map-traversal (e.g., wavefront, random walk, wall bounce) [56] and path-finding (Bug) algorithms [45, 47] for UGV navigation. The equipotential lines labeled as EP* are defined by Eq. 6.
2 Background & Related Work
2.1 Path Planning & Exploration by Mobile Robots
A substantial body of literature has explored various tree search and probabilistic sampling methods for source-to-goal path planning and navigation by UGVs and ASVs [42, 61]. Breadth-First Search (BFS) and Depth-First Search (DFS) are generally adopted for uninformed search, whereas Dijkstra’s algorithm and A* are widely used for informed search [12]. DFS is generally more memory-efficient than BFS and better suited for dynamic obstacles, whereas Dijkstra’s algorithm is ideal when finding the shortest path is essential. Nevertheless, these methods employ exhaustive search, hence can be computationally demanding in complex spaces. The family of A* and D* algorithms [22, 15, 2] improves upon the computational efficiency by incorporating heuristics to reduce unnecessary exploration while guaranteeing the shortest path.
Beyond a discretized grid-like environment, sampling-based algorithms are better suited for continuous high-dimensional spaces [12]. Probabilistic Roadmaps (PRM) [33] and Rapidly-exploring Random Tree (RRT) [39] methods are classical methods that build a tree of random samples from the configuration space, starting from the initial position and incrementally growing toward the goal [74]. PRM is best for multi-query planning in static environments, whereas RRT is fast and suitable for real-time planning but produces suboptimal paths. While some variants, such as RRT*, are asymptotically optimal, although at a higher computational cost [64].
On the other hand, Bug0 [45] and Bug1 [47] algorithms employ a locally optimal strategy of following the line-of-sight and avoiding obstacles along their boundaries. Bug2 [47] optimizes the path using an imaginary m-line (main line) to avoid unnecessary loops. Besides, Alg1 and Alg2 approaches [62] leverage memory of the previous obstacle hit points to search for alternative routes, while DistBug [30] makes it memory efficient by retaining the most recent hit points only. M-Bug [54] minimizes the path length by dynamically adjusting the m-line. Integration of range sensors further allows the robot to sense nearby obstacles and choose a short-cut while maintaining minimum deviation from m-line, as demonstrated in TangentBug [31], WedgeBug [38], and VisBug [46].
In contrast, I-Bug [68] solely relies on signal intensity from the goal to guide navigation. EgressBug [21] offers an escape mechanism, allowing the robot to recover from deadlocks. PointBug [8] introduces the notion of sudden points, i.e., turning points on the outer perimeter of obstacles, which is further improved in E-bug methods [52, 53]. More sophisticated algorithms such as APF-Bug [72] assess ground quality to avoid immobilization in complex terrain obstacles and increase the success rate of reaching the goal.
@@ 2.1 Path Planning & Exploration by Mobile Robots
Figure 3: Our Rover Master platform designed for ClipRover is shown: (a) Back view, showing the single board computer (SBC) and electronics stack including a brushless motor speed controller, a flight controller with IMU, and a power distribution board; (b) Front view, showing the camera for zero-shot navigation; (c) An alternative design, demonstrating the flexibility of this platform, configured with revised wheels, chassis, and additional manipulators for potential field robotics applications.
@@
2.2 Zero-shot Learning & Vision-Language Navigation
@@
Contemporary works on vision-language navigation (VLN) focus on enabling robots to understand language instructions for interactive task planning. VLN algorithms integrates visual information (e.g., recognizing objects, obstacles) and language instructions to plan a trajectory or action sequence for task execution. This is achieved by first generating semantic tags into the SLAM pipeline [25, 17, 20] to generate a 3D map. Sbsequently, the resulting augmented map is used to support online navigation to perform zero-shot semantic tasks [76]. These works primarily focus on adding semantic information to existing mapping and exploration techniques [25, 37], do not participate in the path planning process.
@@
3 Rover Master: System Design
We develop a novel robotic platform “Rover Master” as a low-power and low-cost UGV platform that is portable and scalable for 2D navigation tasks. The major sensors and actuation components are shown in Fig. 3; it includes a monocular RGB camera and a 2D LiDAR for exteroceptive perception. Four independent wheel assemblies are responsible for actuation; each wheel assembly is modular and self-contained, i.e., it consists of a gearbox, a brushless DC motor, and a suspension system; see Fig. 4b. It also includes a pseudo odometer that uses telemetry data from an electronic speed controller (ESC) which drives the motors. Additionally, the driver stack consists of a flight-controller module originally designed for quad-copters that report onboard sensory data to the host computer for planning and navigation. Unlike existing platforms like the TurtleBots [4], RoverMaster is designed with an open and spacious chassis to facilitate easy adaptation to any single board computers (SBCs). In our setup, we configured the platform to: (i) deliver sufficient computational power to handle a general-purpose vision-language model and other computationally intensive tasks; (ii) have enough mechanical stability to hold the camera in an elevated position without excess vibrations even on uneven surfaces; and (iii) support 3-DOF motions (forward/backward, sideways and rotation), whereas most existing UGV systems have only 2-DOF (forward surge and twist rotation).
3.1 Task Specific Design for ClipRover
We customize the Rover Master platform with a configuration that supports real-time VLN capability for 2D environments. An omnidirectional drive system is chosen to account for the moderate surface irregularities. The four wheels are connected to separate brushless DC motors via a planetary gearbox with a reduction ratio of $16$:$1$. This allows the robot to travel at higher speeds and easily overcome moderate obstacles. A throttle limit of $20$% is imposed to ensure operational safety, capping the robot’s maximum speed at approximately $2$ m/s. As shown in Fig. 4b, the gearbox and brushless DC motor are integrated into the wheel hub, optimizing the spatial efficiency and smoothness of the drive system. Moreover, the computing pipeline is powered by an Nvidia Jetson Orin module. A FLIR global shutter RGB camera, lifted approximately $30$ cm above ground, is used for visual perception. This design ensures that the optical center of the lens aligns with the geometric center of the robot. Besides, the heading value from the flight controller’s IMU facilitates $360^{\circ}$ scanning capabilities. A 2D LiDAR is mounted at the front of the upper deck as is shown in Fig. 3b. The LiDAR is only used for safety and visualization purposes and does not influence VLN’s decision-making. Further details regarding the LiDAR’s utility are explained in Sec. 5.4.
@@ 3.2 Features and Capabilities
One unique advantage of our Rover Master system design is the compactness of its wheels and motor assembly. The four-wheel independent suspension extends its application to uneven surfaces; when using the Discovery Wheel (TPU) shown in Fig. 3c and Fig. 4b, it performs well on grasslands, rocky pavements, and even sand (with reduced maneuverability). The suspension system can also be configured to filter out surface bumps and provide better stability for onboard sensors.
The Rover Master platform can be configured to switch between different drive types (omnidirectional or differential), chassis sizes (with parameterized CAD design), and task-specific actuation and sensory add-ons (cameras, LiDARs, manipulators for both indoor and field robotics applications). The chassis plates are designed to house various additional sensors and actuators according to tasks. A comparison of Rover Master with two widely used TurtleBot UGV variants is presented in Table 1.
@@
4 ClipRover: Navigation Pipeline
Recent advancements in vision-language models (VLMs) have demonstrated their potential for spatial reasoning and awareness capabilities [10, 14]. Different from traditional approaches, this study explores the feasibility of incorporating VLMs directly into a robot’s autonomy pipeline to facilitate real-time exploration and target identification. To this end, we introduce a modular pipeline that integrates stages for perception, planning, and navigation. The core computational components of the proposed ClipRover navigation pipeline are depicted in Fig. 5.
The proposed pipeline comprises three main stages. The frontend processes raw input frames, divides them into tiles, and encodes these tiles into embeddings—numerical vectors representing semantic meanings; In our implementation, the CLIP vision encoder is employed as the frontend. Then, middlewares components take the visual embeddings generated by the frontend as input and produce scores that carry specific semantic interpretations. These scores are typically derived by correlating the input embeddings with the middleware’s internal database. The meanings of these scores can vary depending on the application’s requirements. In this study, the middleware generates three types of scores: navigability, familiarity, and target confidence. Lastly, at the backend, those scores from middlewares are used to make motion decisions. It is designed to be adaptable, allowing the integration of different algorithms tailored to various applications and environments. For this study, a minimal backend was implemented with three operational modes: basic navigation, look-around, and target lock. These modes are dynamically activated or deactivated based on the provided scores.

### v2

# VL-Explore: Zero-shot Vision-Language Exploration and Target Discovery by Mobile Robots

@@ Abstract
Vision-language navigation (VLN) has emerged as a promising paradigm, enabling mobile robots to perform zero-shot inference and execute tasks without specific pre-programming. However, current systems often separate map exploration and path planning, with exploration relying on inefficient algorithms due to limited (partially observed) environmental information. In this paper, we present a novel navigation pipeline named “VL-Explore” for simultaneous exploration and target discovery in unknown environments, leveraging the capabilities of a vision-language model named CLIP. Our approach requires only monocular vision and operates without any prior map or knowledge about the target. For comprehensive evaluations, we designed a functional prototype of a UGV (unmanned ground vehicle) system named “Open Rover”, a customized platform for general-purpose VLN tasks. We integrated and deployed the VL-Explore  pipeline on Open Rover  to evaluate its throughput, obstacle avoidance capability, and trajectory performance across various real-world scenarios. Experimental results demonstrate that VL-Explore  consistently outperforms traditional map-traversal algorithms and achieves performance comparable to path-planning methods that depend on prior map and target knowledge. Notably, VL-Explore  offers real-time active navigation without requiring pre-captured candidate images or pre-built node graphs, addressing key limitations of existing VLN pipelines.
@@ 1 Introduction
Autonomous robots face considerable challenges in exploring and discovering targets within unknown environments without a prior map, typically requiring a dedicated mapping phase before initiating path planning strategies [12, 23, 8]. Recent vision-language (VL) approaches have begun to address this by incorporating long-term memory architectures, a key factor driving performance gains in navigation tasks [6]. However, methods that rely on pre-constructed maps or pre-existing environmental knowledge often exhibit poor adaptability to dynamic or evolving conditions, necessitating frequent reinitialization or bespoke algorithmic interventions [15, 1]. These limitations are especially pronounced in critical real-world applications—such as search-and-rescue, environmental monitoring, and warehouse exploration—where adaptability and robustness are essential [20, 40, 26]. Despite the importance of efficient, zero-shot exploration in such scenarios, current VLN systems have yet to adequately address this capability.
Figure 1: (a) Existing VLN systems generally use VLMs as a complementary add-on; (b) The proposed VL-Explore pipeline is designed with a general-purpose VLM at its core, eliminating the need for multimodal sensing and extra processing steps.
Figure 2: Illustration of an ongoing exploration and target discovery task by the proposed VL-Explore  system. Left: A synthesized image representing the robot’s perspective from its onboard camera; the cone illustrates its field-of-view, divided into six numbered tiles for detailed analysis. Right: The vision-language perception of each tile is processed by a novel correlation middleware; The motion mixer engine is currently in ‘target lock’ mode, prioritizing the target (a teddy bear) in tile-5 over other navigable regions (e.g., tile-6) due to the higher precedence assigned to the target.
When initializing inside an unexplored environment, many traditional methods use a fixed policy that often gets stuck in local minima, leading to inefficient or incomplete exploration [21]. To address this, contemporary VLN systems use high-sensing modalities [12, 6] such as multiple cameras, depth camera or LiDAR to help with first-time exploration. For instance, a 2D LiDAR is commonly used for obstacle avoidance and mapping [38, 35], while a separate camera is used for scene classification and target detection [22]. Additionally, the initial map traversal is not feasible for partially observable and dynamic environments, requiring frequent re-initialization [16, 5]. To this end, the efficiency and robustness of target discovery by zero-shot exploration remain largely under-explored.
In this paper, we present “VL-Explore”, a novel navigation framework for zero-shot exploration and target discovery by unmanned ground vehicles (UGVs) in unknown environments. VL-Explore extends the spatial context awareness capabilities of general-purpose VLMs [7] to guide 2D robotic exploration and target discovery without requiring access to a prior map. It addresses the limitation that existing VLN systems predominantly use a VLM as a complementary add-on for policy projection, outside the navigation subsystem; see Fig. 1. Instead, VL-Explore  uses a general-purpose VLM as its navigation core, with the entire system operating within it. It also eliminates the need for any additional sensing modalities and map building. As illustrated in Fig. 1b, simplifying the system architecture in this way significantly reduces the system complexity and cost, making it more scalable and suitable for resource-constrained systems.
VL-Explore eliminates the contemporary VLM system’s reliance on pre-built knowledge graphs and candidate image libraries. Specifically, SOTA systems such as Clip-Nav [9], Clip on Wheels [10], and GCN [23] require pre-captured images of candidate scenes, target locations, and/or pre-built knowledge graphs of the navigable space as prior [8]. Due to this complexity, they are generally tested in 2D simulations or overly simplistic environments. Other contemporary methods, such as SEEK [12], require high-sensing modalities involving multiple cameras, LiDAR, and onboard SLAM pipelines for navigation. In VL-Explore, we formulate the problem from a different perspective. Instead of relying on a separate traditional mapping and path-planning backbone, our proposed architecture uses the visual embeddings encoded by a foundational VLM as the only input to the system. High-level functionalities such as obstacle avoidance, path planning, and target identification are built upon this foundation for active visual perception.
To validate the proposed framework in real-world environments, we develop a UGV platform that meets the computational demands of VLMs while offering superior maneuverability and mechanical stability during the task. We deployed the CLIP model [36, 18] in our proposed architecture and performed extensive tests in a real-world environment. The results demonstrate that VL-Explore consistently outperforms map-traversal algorithms and achieves performance comparable to path-finding methods, despite the latter relying on prior knowledge of the map and target location. Moreover, VL-Explore achieves significantly shorter trajectory lengths, making it more efficient than map-traversal algorithms. In comparison with state-of-the-art systems, VL-Explore achieves comparable or even superior performance to systems with higher sensor modalities or even with long-term memory. Notably, VL-Explore offers such performance margins without additional sensor modalities, also requiring no separate dedicated map-building or localization modules.
Overall, we make the following contributions in this paper:
@@
    We propose a model-centric navigation pipeline “VL-Explore” for simultaneous exploration and target discovery by UGVs in unknown environments. It requires no additional sensor modalities other than RGB feed for zero-shot active navigation.
@@
    We design a novel evaluation metric “Entropy Preserving Score” (EPS) to measure the exploration efficiency, improving the consistency of score across different experimental environments hence providing better comparability across different systems.
@@
    We perform extensive experiments with the proposed system and compare its performance against other systems, demonstrating the effectiveness of the proposed system in real-world environments.
@@
Figure 3: Our Open Rover  platform designed for VL-Explore  is shown: (a) Back view, showing the single board computer (SBC) and electronics stack including a brushless motor speed controller, a flight controller with IMU, and a power distribution board; (b) Front view, showing the camera for zero-shot navigation; (c) An alternative design, demonstrating the flexibility of this platform, configured with revised wheels, chassis, and additional manipulators for potential field robotics applications.
@@
2 Related Work: Zero-shot Learning and Vision-Language Navigation
@@
Contemporary works on vision-language navigation (VLN) focus on enabling robots to understand language instructions for interactive task planning. VLN algorithms integrate visual information (e.g., recognizing objects, obstacles) and language instructions to plan a trajectory or action sequence for task execution. This is achieved by first generating semantic tags into the SLAM pipeline [17, 10, 14] to generate a 3D map. Subsequently, the resulting augmented map is used to support online navigation to perform zero-shot semantic tasks [42]. These works primarily focus on adding semantic information to existing mapping and exploration techniques [17, 24], but do not participate in the path planning process.
@@
3 Open Rover: Platform Design
We develop a novel robotic platform “Open Rover”. The platform is designed to be flexible with drive types, chassis sizes, sensor and actuation add-ons, so it can adapt to the needs of prospective research. The major sensors and actuation components are shown in Fig. 3; it includes a monocular RGB camera and a 2D LiDAR for exteroceptive perception. Four independent wheel assemblies are responsible for actuation; each wheel assembly is modular and self-contained, i.e., it consists of a gearbox, a brushless DC motor, and a suspension system; see Fig. 4b. It also includes a pseudo-odometer that uses telemetry data from an electronic speed controller (ESC), which drives the motors. Additionally, the driver stack consists of a flight-controller module originally designed for quadcopters that reports onboard sensory data to the host computer for planning and navigation.
Unlike existing platforms like the TurtleBots [2], Open Roveris designed with an open and spacious chassis to facilitate easy adaptation to any single board computer (SBC). In our setup, we configured the platform to: (i) deliver sufficient computational power to handle a general-purpose vision-language model and other computationally intensive tasks; (ii) have enough mechanical stability to hold the camera in an elevated position without excess vibrations even on uneven surfaces; and (iii) support 3-DOF motions (forward/backward, sideways and rotation), whereas most existing UGV systems have only 2-DOF: forward surge and twist rotation.
3.1 Task Specific Design for VL-Explore
We customize the Open Rover  platform with a configuration that supports real-time VLN capability for 2D environments. An omnidirectional drive system is chosen to account for the moderate surface irregularities. The four wheels are connected to separate brushless DC motors via a planetary gearbox with a reduction ratio of $16$:$1$. This allows the robot to travel at higher speeds and easily overcome moderate obstacles. A throttle limit of $20$% is imposed to ensure operational safety, capping the robot’s maximum speed at approximately $2$ meters per second.
As shown in Fig. 4b, the gearbox and brushless DC motor are integrated into the wheel hub, optimizing the spatial efficiency and smoothness of the drive system. Moreover, the computing pipeline is powered by an Nvidia Jetson Orin module. A FLIR global shutter RGB camera, lifted approximately $30$ cm above ground, is used for visual perception. This design ensures that the optical center of the lens aligns with the geometric center of the robot. Besides, the heading value from the flight controller’s IMU facilitates $360^{\circ}$ scanning capabilities. A 2D LiDAR is mounted at the front of the upper deck as is shown in Fig. 3b. The LiDAR is only used for safety and visualization purposes and does not influence VLN’s decision-making. Further details regarding the LiDAR’s utility are explained in Sec.4 of the supplementary materials.
@@ 3.2 Features and Capabilities
One unique advantage of our Open Rover  system design is the compactness of its wheels and motor assembly. The four-wheel independent suspension extends its application to uneven surfaces; when using the Discovery Wheel (TPU) shown in Fig. 3c and Fig. 4b, it performs well on grasslands, rocky pavements, and even sand (with reduced maneuverability). The suspension system can also be configured to filter out surface bumps and provide better stability for onboard sensors.
@@
The Open Rover  platform can be configured to switch between different drive types (omnidirectional or differential), chassis sizes (with parameterized CAD design), and task-specific actuation and sensory add-ons (cameras, LiDARs, manipulators for both indoor and field robotics applications). The chassis plates are designed to house various additional sensors and actuators according to tasks. A comparison of Open Rover  with two widely used TurtleBot UGV variants is presented in Table 1.
4 VL-Explore: Navigation Pipeline
Recent advancements in VLMs have demonstrated their potential for spatial reasoning and awareness capabilities [7, 9]. Different from traditional approaches, this study explores the feasibility of incorporating VLMs directly into a robot’s autonomy pipeline to facilitate real-time exploration and target identification. To this end, we introduce a modular pipeline that integrates stages for perception, planning, and navigation. The core computational components of the proposed VL-Explore   navigation pipeline are depicted in Fig. 5.
The proposed pipeline comprises three main stages. The frontend processes raw input frames, divides them into tiles, and encodes these tiles into embeddings—numerical vectors representing semantic meanings; In our implimentation, the CLIP vision encoder is employed as the frontend. Then, middlewares take the visual embeddings generated by the frontend as input and produce scores that carry specific semantic interpretations. These scores are typically derived by correlating the input embeddings with the middleware’s internal database. The meanings of these scores can vary depending on the application’s requirements. In this study, the middleware generates three types of scores: navigability, familiarity, and target confidence.
Lastly, at the backend, those scores from middlewares are used to make motion decisions. It is designed to be adaptable, allowing the integration of different algorithms tailored to various applications and environments. For this study, a minimal backend was implemented with three operational modes: basic navigation, look-around, and target lock. These modes are dynamically activated or deactivated based on the provided scores.
Notably, the proposed VL-Explore   pipeline is agnostic to CLIP or any other VL backbone. OpenCLIP is used in our experiments given its open-source status and 3-4 years of maturity in various research fields, ensuring reproducibility. While a more recent multi-modal VLM could potentially enhance the performance of the suggested pipeline, the primary focus of this work is to validate the feasibility of integrating VLMs into a real-time navigation system. Thus, we chose to utilize a well-established model for better reproducibility.